# Logarithmic regret in the ergodic Avellaneda–Stoikov market making model

### v1

@@ 3. Numerical experiments
In this section, we numerically simulate the ergodic market making model (7) and the achieved regret of Algorithm 1.
Let us consider the following parameters in the simulation: $\lambda^{\pm}=1/s$, $\kappa^{\pm}=10\$^{-1}$, $\sigma=1.0s^{-1/2}\$$, $S_{0}=\$10$, $\bar{q}=30$, $\underline{q}=-30$ and $\phi=\$1\times 10^{-5}$. In the learning part, we set $\kappa^{\star}=10\$^{-1}$, $\underline{K}=1\$^{-1}$ and $\bar{K}=100\$^{-1}$.
By Theorem 2, we can determine the square matrix $\bm{A}$ and the largest eigenvalue of $\bm{A}$ is $\lambda_{max}=0.7297$. Then we have $\gamma=0.07297$ by using Theorem 6. Figure 1 (left panel) plots the asymptotic behaviors of $v(0,q;T)/T$ as $T$ increases, which $v(t,q;T)$ is the value function in the discounted finite-time-horizon setting with the discount factor $r=0$ given by (12). We can see that, for any initial $q\in\Omega^{Q}$, the value $v(0,q;T)/T\rightarrow\gamma=\lambda_{max}/\kappa$ as $T\rightarrow+\infty$ as stated in Theorem 5.
Figure 1. Left: The asymptotic behaviors of $v(0,q;T)/T$ as $T$ increases. Middle: Several possible solutions $\hat{v}(q)$ to the ergodic HJB equation (16). Right: The unique optimal control $\delta^{\star}$.
@@
Now we proceed to simulate the proposed learning algorithm for the price sensitivity of market participants parameter $\kappa$ and its achieved regret. Figure 3 (left panel) plots the learning error $\left|\kappa_{t}-\kappa^{\star}\right|$ of Algorithm 1 in the log-log scale. Initially, the learning rate is slow due to the limited number of Bernoulli signals. As time increases, the estimated $\kappa_{t}$ quickly converges to the true parameter $\kappa^{\star}$. Moreover, the right panel of Figure 3, which is the Monte-Carlo simulation of the achieved (expected) regret of Algorithm 1, demonstrates that the algorithm regret grows sublinearly and is bounded by order $\mathcal{O}(\ln^{2}T)$. The complete code can be found at https://github.com/Galen-Cao/MM_parmater_learning.
@@
Figure 3. Left: Log-log plot of the learning error $\left|\kappa_{t}-\kappa^{\star}\right|$ over time under Algorithm 1. Right: The regret over time for Algorithm 1.
@@ 4. Conclusion
In this paper, we introduced and analysed the ergodic formulation of the Avellaneda–Stoikov market making model. We established explicit solutions to the ergodic Hamilton–Jacobi–Bellman (HJB) equation and thus derived the optimal ergodic Markov controls. We’ve further shown that under the ergodic optimal control there is a unique invariant distribution for the market maker’s inventory and that any initial distribution converges exponentially fast to the equilibrium one. This allowed us to establish the regret upper bound of $\mathcal{O}(\ln^{2}T)$ for learning the unknown price sensitivity of liquidity takers $\kappa^{\star}$. Our work extends the known results on the market making model by providing a rigorous analysis of the ergodic setting and offering a robust solution for parameter learning. The numerical experiments further validate the theoretical results, confirming the robustness of the proposed algorithm.
@@ A.1. Proof of Theorem 1
$\begin{split}\lambda^{+}e^{-\kappa^{+}\delta^{+}}(p_{1}+\delta^{+})\bm{1}_{q>\underline{q}}+&\lambda^{-}e^{-\kappa^{-}\delta^{-}}(p_{2}+\delta^{-})\bm{1}_{q<\bar{q}}-\phi q^{2}\,>\,\\ &\lambda^{+}e^{-\kappa^{+}\delta^{+}}(p^{\prime}_{1}+\delta^{+})\bm{1}_{q>\underline{q}}+\lambda^{-}e^{-\kappa^{-}\delta^{-}}(p_{2}+\delta^{-})\bm{1}_{q<\bar{q}}-\phi q^{2}\end{split}$
@@
To prove the existence of the unique solution $u$ to the HJB equation (13) on $(-\infty,T]$, we use Lemma 35 (4) to prove the locally Lipschitz of the ODE and apply [27, Theorem 3.3].
@@ A.2. Proof of Theorem 3
To see that, $\forall q\in\Omega^{Q}$ and $t\in\mathbb{R}^{+}$, $v_{r}(q)=\lim_{T\rightarrow+\infty}v_{r}(t,q;T)$, we apply [27, Proposition 4.1] by using the running reward function $f$ is bounded. Moreover, by a standard verification argument, we know that $v_{r}$ is the solution to the HJB equation (15), hence the result of Theorem 3. ∎
@@ A.3. Proof of Theorem 4
By using the fact that the running reward function $f$ is bounded (33) and (34), we can apply [27, Lemma 4.3(1)] to get statement (i) of the lemma.
To prove statement (ii), we first define the stopping time $\tau(q,\hat{q})$ under a control $\delta^{\pm}\in\mathcal{A}$ as
$\tau:=\inf\{t\left.\right|Q_{t}^{\delta^{\pm},0,q}=\hat{q}\}\,.$
With the fact that, for $\forall q,\hat{q}\in\Omega^{Q}$, the transition probability for $\hat{q}$ from $q$, $\bm{P}(q,\hat{q})>0$, we have $\mathbb{E}[Q_{\tau}]<+\infty$. See more detailed discussion in Appendix A.10 for the transition probability matrix.
Let us consider $\delta^{\pm,\epsilon}\in\mathcal{A}[0,\tau]$ with $\epsilon>0$ such that
$v_{r}(q)-\epsilon\leq\mathbb{E}_{q}\Big{[}\int_{0}^{\tau}e^{-rt}f(Q_{t};\delta^{\pm,\epsilon})\,\mathrm{d}t+e^{-r\tau}v_{r}(\hat{q})\Big{]}\,.$
@@
$\begin{split}v_{r}(q)-\epsilon-\frac{\underline{C}}{r}&\leq\mathbb{E}_{q}\Big{[}\int_{0}^{\tau}e^{-rt}\big{(}f(Q_{t};\delta^{\pm})-\underline{C}\big{)}\,\mathrm{d}t+e^{-r\tau}\big{(}v_{r}(\hat{q})-\frac{\underline{C}}{r}\big{)}\Big{]}\\ &\leq\mathbb{E}\Big{[}\int_{0}^{\tau}e^{-rt}\big{(}f(Q_{t};\delta^{\pm})-\underline{C}\big{)}\,\mathrm{d}t\Big{]}+\mathbb{E}\big{[}e^{-r\tau}\big{]}\big{(}v_{r}(\hat{q})-\frac{\underline{C}}{r}\big{)}\\ &\leq(\bar{C}-\underline{C})\mathbb{E}[\tau]+v_{r}(\hat{q})-\frac{\underline{C}}{r}\,.\end{split}$
@@
$\begin{split}\frac{v_{r}(q)-v_{r}(\hat{q})}{|q-\hat{q}|}&\leq\frac{1}{|q-\hat{q}|}\Big{(}(\bar{C}-\underline{C})\mathbb{E}[\tau]+\epsilon\Big{)}\leq(\bar{C}-\underline{C})\mathbb{E}[\tau]+\epsilon,\quad q,\hat{q}\in\Omega^{Q},q\neq\hat{q}\,.\end{split}$
Since $\mathbb{E}[\tau]<+\infty$ and $q\in\Omega^{Q}$, by letting $\epsilon\rightarrow 0$, we conclude that $v_{r}(q)-v_{r}(\hat{q})$ is bounded from above. By simply changing the order of $q$ and $\hat{q}$, we conclude the lower boundedness. Hence, we can find $C_{2}\in\mathbb{R}$ such that
@@
In the proof we follow the ideas from [27, Proposition 4.6, 4.7]. As $\left|rv_{r}(q)\right|\leq C_{1}$ and $\left|v_{r}(q)-v_{r}(0)\right|\leq C_{2}\bar{q}=C^{\prime}_{2},\,\forall q\in\Omega^{Q}$ by Lemma 36, we can consider a sequence $(r_{n})_{n\in\mathbb{N}}$ converging towards 00 such that the sequences $\big{(}r_{n}v_{r_{n}}(q)\big{)}_{n\in\mathbb{N}}$ and $\big{(}v_{r_{n}}(q)-v_{r_{n}}(0)\big{)}_{n\in\mathbb{N}}$ are convergent for $q\in\Omega^{Q}$. Let $\hat{\gamma}(q)$ denote the limit of the sequence $\big{(}r_{n}v_{r_{n}}(q)\big{)}_{n\in\mathbb{N}}$, we have
@@
Let $z(q)=\hat{w}(q)-\hat{v}(q)$. Since the domain for $z(q)$ is bounded, we know that the supremum and infimum exist. Let us denote $\bar{z}=\sup_{q\in\Omega^{Q}}z(q)$, $\underline{z}=\inf_{q\in\Omega^{Q}}z(q)$ and $\epsilon=\frac{\hat{\gamma}-\eta}{\bar{z}-\underline{z}+1}$.
Let us first assume $\hat{\gamma}>\eta$ and prove that $\hat{\gamma}\leq\eta$ by contradiction. By the definition of $\epsilon$, we have, for $\forall q\in\Omega^{Q}$,
$\begin{split}0\leq\epsilon&(\bar{z}-z(q)+1)\leq\hat{\gamma}-\eta\\ &=H\Big{(}q,(\hat{v}(q^{\prime})-\hat{v}(q))_{q^{\prime}\in\{q-1,q+1\}}\Big{)}-H\Big{(}q,(\hat{w}(q^{\prime})-\hat{w}(q))_{q^{\prime}\in\{q-1,q+1\}}\Big{)}\,.\end{split}$
@@
$\begin{split}-\epsilon\hat{w}(q)+H\Big{(}q,(\hat{w}(q^{\prime})-&\hat{w}(q))_{q^{\prime}\in\{q-1,q+1\}}\Big{)}\leq\\ &-\epsilon(\hat{v}(q)+\bar{z}+1)+H\Big{(}q,(\hat{v}(q^{\prime})-\hat{v}(q))_{q^{\prime}\in\{q-1,q+1\}}\Big{)}\,.\end{split}$
By using the comparison principle, see [27, Lemma 4.4], we know that $\hat{v}(q)+\bar{z}+1\leq\hat{w}(q)$, for $\forall q\in\Omega^{Q}$, which indicates a contradiction with the definition of $\bar{z}$. Hence, we have $\hat{\gamma}\leq\eta$. By simply changing the order of $\hat{\gamma}$ and $\eta$, we can obtain that $\hat{\gamma}\geq\eta$. Therefore, we conclude that $\hat{\gamma}=\eta$, i.e. $\hat{\gamma}$ is independent of the sequence $(r_{n})_{n\in\mathbb{N}}$. ∎

### v2

@@ 3. Numerical experiments
3.1. Ergodic control and regret of Algorithm 1
In this section, we numerically simulate the ergodic market making model (7) and the achieved regret of Algorithm 1. Code used to produce results in this section is available at https://github.com/Galen-Cao/MM_parmater_learning.
Let us consider the following parameters in the simulation: $\lambda^{\pm}=1/s$, $\kappa^{\pm}=10\$^{-1}$, $\sigma=1.0s^{-1/2}\$$, $S_{0}=\$10$, $\bar{q}=30$, $\underline{q}=-30$ and $\phi=\$1\times 10^{-5}$.
By Theorem 2, we can determine the square matrix $\bm{A}$ and the largest eigenvalue of $\bm{A}$ is $\lambda_{max}=0.7297$. Then we have $\gamma=0.07297$ by using Theorem 6. Figure 1 (left panel) plots the asymptotic behaviours of $v(0,q;T)/T$ as $T$ increases, which $v(t,q;T)$ is the value function in the discounted finite-time-horizon setting with the discount factor $r=0$ given by (12). We can see that, for any initial $q\in\Omega^{Q}$, the value $v(0,q;T)/T\rightarrow\gamma=\lambda_{max}/\kappa$ as $T\rightarrow+\infty$ as stated in Theorem 5.
Figure 1. Left: The asymptotic behaviours of $v(0,q;T)/T$ as $T$ increases. Middle: Several possible solutions $\hat{v}(q)$ to the ergodic HJB equation (16). Right: The unique optimal control $\delta^{\ast}$.
@@ 3.1. Ergodic control and regret of Algorithm 1
Figure 3. Top left: Log-log plot of the learning error $\left|\kappa_{t}-\kappa^{\ast}\right|$ over time under Algorithm 1. Top right: The regret of Algorithm 1 over time. Bottom: The learning error and regret over a larger time horizon.
Now we proceed to simulate learning and the regret of Algorithm 1. The results are in Figure 3. The experimental parameters are set as $\lambda^{\pm}=0.4/s$, $\kappa^{\pm\ast}=10\$^{-1}$, $\sigma=0.01s^{-1/2}\$$, $\bar{q}=30$, $\underline{q}=-30$, $\phi=\$1\times 10^{-6}$, $\underline{K}=1\$^{-1}$ and $\bar{K}=100\$^{-1}$. We used $1000$ simulation scenarios, with time horizon $T=1000$ seconds. We include two plots, one with $T=100$ seconds and one with $T=1000$ seconds. The left panels show the learning error $\left|\kappa_{t}-\kappa^{\ast}\right|$ in the log-log scale. Initially, the error decays slowly due to the limited number of Bernoulli signals. However, as time increases, the estimate $\kappa_{t}$ rapidly converges to the true value $\kappa^{\ast}$, demonstrating the algorithm’s consistency. The right panels illustrate the Monte Carlo simulation of the regret achieved by Algorithm 1. Under both time horizons, the regret shows sublinear growth and is bounded by order $\mathcal{O}(\ln^{2}T)$. Curve fitting analysis further confirms that, especially at the longer time scale, the curve of order $\mathcal{O}(\ln^{2}T)$ provides a better fit than that of order $\mathcal{O}(\ln T)$, which supports our theoretical regret analysis.
Furthermore, we compare Algorithm 1 with a myopic benchmark strategy that always posts at $\frac{1}{\kappa_{t}}$, where $\kappa_{t}$ is the current estimate of $\kappa^{\ast}$ at time $t$. As shown in Figure 4, while the myopic strategy is still able to learn the true parameter over time, the corresponding regret grows linearly with time. In contrast, Algorithm 1 achieves sublinear regret, implying the advantages of employing the “optimal” (computed from the estimate $\kappa_{t}$) policies in reducing regret. Note that this is the regret of a risk-averse market maker with risk aversion parameter $\phi=\$10^{-6}$.
Figure 4. Performance comparison between Algorithm 1 and a myopic benchmark strategy that posts at $1/\kappa_{i}$ using the current estimate. Left: Log-log plot of the estimation error $|\kappa_{t}-\kappa^{\ast}|$. Right: Regret over time.
3.2. Non-stationary market
Financial markets are typically non-stationary. To handle the non-stationarity of $\kappa$, we incorporate two classical techniques in our learning algorithm: a sliding-window (SW) approach and an exponential-weighted-moving-average (EWMA) approach.
The sliding window (SW) method is as follows. The index set of recent data is defined as $\mathcal{I}_{i}:=\{j\leq i\mid t_{i}-t_{j}\leq w\}$. We then obtain $\kappa_{i}^{(w)}$ by numerically solving $\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{i}^{(w)}(\kappa)=0$, where the expression for the derivative $\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{i}^{(w)}$ is given by (28), but computed using only the data points indexed by $\mathcal{I}_{i}$, i.e. from $j=\inf\mathcal{I}_{i}$ to $j=i$. Otherwise, the algorithm is the same as Algorithm 1.
The exponential-weighted-moving-average (EWMA) method uses the following log-likelihood
$\ell_{N}^{EWMA}(\kappa)=\sum_{n=1}^{N}e^{-\alpha(t_{N}-t_{n})}\left(-\kappa\delta_{n}Y_{n}+(1-Y_{n})\log(1-e^{-\kappa\delta_{n}})\right)\,,$
where $\alpha$ is the weighting parameter. This is then regularised as in (27) where $\ell_{N}$ is replaced by $\ell_{N}^{EWMA}$. The algorithm incorporating the EWMA method simply replaces the log-likelihood function in Algorithm 1 with regularisation of $\ell_{N}^{EWMA}$.
Figure 5 illustrates the performance of the learning algorithms in a non-stationary market environment. We use the same parameters as those used for Figure 3, except in this non-stationary setting, the true value of $\kappa$ changes every $50$ seconds, following the sequence $[20,30,10,40,25]$. The SW algorithm employs a sliding window of $30$ seconds, while the EWMA algorithm sets $\alpha=0.1$. The left panel shows how the estimated $\kappa$ (green and orange curves) tracks the true, piecewise constant $\kappa$ (blue dashed line) over time using each method. We observe that after each shift in the true value, the estimate gradually converge to the new value, with a short delay in both methods. The right panel presents the regrets of the two methods over time. As expected, the regret grows approximately at an order of $\ln^{2}T$ in each regime where $\kappa$ is fixed. However, each change in $\kappa$ introduces a noticeable increase in the regret due to the lag in adaptation. Once the estimates converge to the new value, the growth of regret slows down again. Even though with our choice of window size and $\alpha$ the SW method achieves lower regret than the EWMA method this does not imply that the SW method is better; we expect there will be a value of $\alpha$ where the EWMA method achieves the same regret.
Figure 5. Learning $\kappa$ in the non-stationary market
3.3. The dependence of the regret bound on model parameters
In this section, we implement numerical experiments to quantify the dependence of the regret constant $C_{1}$ given in Theorem 17 on the key model parameters, $\phi,\lambda^{\pm},\bar{K},\underline{K}$ and $\kappa_{0}-\kappa^{\ast}$. Figure 6 (left panel) presents how $C_{1}$ varies with $\phi$ and $\lambda$, where we set $\lambda^{\pm}=\lambda$ in the simulation. We used $500$ scenarios and $T=100$ seconds with the random seed fixed for each parameter combination. We observe that $C_{1}$ increases as both $\phi$ and $\lambda$ increase. This is expected as larger $\phi$ implies a higher penalty for the holding inventory, while a larger $\lambda$ corresponds to a higher frequency of incoming market orders. The right panel shows the dependence of $C_{1}$ on $\bar{K}$, $\underline{K}:=\frac{1}{\bar{K}}$ and on $\kappa_{0}-\kappa^{\ast}$. As shown in the figure, a higher $\bar{K}$ or higher $\kappa_{0}-\kappa^{\ast}$ leads to a larger constant $C_{1}$ in the asymptotic expression for regret.
Figure 6. Dependence of regret bound constant $C_{1}$ on model parameters
@@ 4. Conclusion
In this paper, we introduced and analysed the ergodic formulation of the Avellaneda–Stoikov market making model. We established explicit solutions to the ergodic Hamilton–Jacobi–Bellman (HJB) equation and thus derived the optimal ergodic Markov controls. We’ve further shown that under the ergodic optimal control there is a unique invariant distribution for the market maker’s inventory and that any initial distribution converges exponentially fast to the equilibrium one. This allowed us to establish the regret upper bound of $\mathcal{O}(\ln^{2}T)$ for learning the unknown price sensitivity of liquidity takers $\kappa^{\ast}$. Our work extends the known results on the market making model by providing a rigorous analysis of the ergodic setting and offering a robust solution for parameter learning. The numerical experiments further validate the theoretical results, confirming the robustness of the proposed algorithm.
@@ A.1. Proof of Theorem 1
$\begin{split}\lambda^{+}e^{-\kappa^{+}\delta^{+}}(p_{1}+\delta^{+})\mathds{1}_{q>\underline{q}}+&\lambda^{-}e^{-\kappa^{-}\delta^{-}}(p_{2}+\delta^{-})\mathds{1}_{q<\bar{q}}-\phi q^{2}\,>\,\\ &\lambda^{+}e^{-\kappa^{+}\delta^{+}}(p^{\prime}_{1}+\delta^{+})\mathds{1}_{q>\underline{q}}+\lambda^{-}e^{-\kappa^{-}\delta^{-}}(p_{2}+\delta^{-})\mathds{1}_{q<\bar{q}}-\phi q^{2}\end{split}$
@@
To prove the existence of the unique solution $u$ to the HJB equation (13) on $(-\infty,T]$, we use Lemma 37 (4) to prove the locally Lipschitz of the ODE and apply [26, Theorem 3.3].
@@ A.2. Proof of Theorem 3
To see that, $\forall q\in\Omega^{Q}$ and $t\in\mathbb{R}^{+}$, $v_{r}(q)=\lim_{T\rightarrow+\infty}v_{r}(t,q;T)$, we apply [26, Proposition 4.1] by using the running reward function $f$ is bounded. Moreover, by a standard verification argument, we know that $v_{r}$ is the solution to the HJB equation (15), hence the result of Theorem 3. ∎
@@ A.3. Proof of Theorem 4
By using the fact that the running reward function $f$ is bounded (37) and (38), we can apply [26, Lemma 4.3(1)] to get statement (i) of the lemma.
To prove statement (ii), we first define the stopping time $\tau(q,\hat{q})$ for the process $Q_{t}$ under a control $\delta^{\pm}\in\mathcal{A}$ with initial condition $Q_{0}=q$ as
$\tau:=\inf\{t\left.\right|Q_{t}^{\delta^{\pm},q}=\hat{q}\}\,.$
Since the dynamics of $Q_{t}^{\delta^{\pm},q}$ can be equivalently represented by a continuous-time Markov chain that is irreducible and recurrent (see the detailed discussion in Section A.15, with $\psi$ replaced by $\delta^{\pm}$), it follows that $\mathbb{E}[\tau]<+\infty$.
Let us consider $\delta^{\pm,\varepsilon}\in\mathcal{A}[0,\tau]$ with $\varepsilon>0$ such that
$v_{r}(q)-\varepsilon\leq\mathbb{E}_{q}\Big{[}\int_{0}^{\tau}e^{-rt}f(Q_{t};\delta^{\pm,\varepsilon})\,\mathrm{d}t+e^{-r\tau}v_{r}(\hat{q})\Big{]}\,.$
@@
$\begin{split}v_{r}(q)-\varepsilon-\frac{\underline{C}}{r}&\leq\mathbb{E}_{q}\Big{[}\int_{0}^{\tau}e^{-rt}\big{(}f(Q_{t};\delta^{\pm})-\underline{C}\big{)}\,\mathrm{d}t+e^{-r\tau}\big{(}v_{r}(\hat{q})-\frac{\underline{C}}{r}\big{)}\Big{]}\\ &\leq\mathbb{E}\Big{[}\int_{0}^{\tau}e^{-rt}\big{(}f(Q_{t};\delta^{\pm})-\underline{C}\big{)}\,\mathrm{d}t\Big{]}+\mathbb{E}\big{[}e^{-r\tau}\big{]}\big{(}v_{r}(\hat{q})-\frac{\underline{C}}{r}\big{)}\\ &\leq(\bar{C}-\underline{C})\mathbb{E}[\tau]+v_{r}(\hat{q})-\frac{\underline{C}}{r}\,.\end{split}$
@@
$\begin{split}\frac{v_{r}(q)-v_{r}(\hat{q})}{|q-\hat{q}|}&\leq\frac{1}{|q-\hat{q}|}\Big{(}(\bar{C}-\underline{C})\mathbb{E}[\tau]+\varepsilon\Big{)}\leq(\bar{C}-\underline{C})\mathbb{E}[\tau]+\varepsilon,\quad q,\hat{q}\in\Omega^{Q},q\neq\hat{q}\,.\end{split}$
Since $\mathbb{E}[\tau]<+\infty$ and $q\in\Omega^{Q}$, by letting $\varepsilon\rightarrow 0$, we conclude that $v_{r}(q)-v_{r}(\hat{q})$ is bounded from above. By simply changing the order of $q$ and $\hat{q}$, we conclude the lower boundedness. Hence, we can find $C_{2}\in\mathbb{R}$ such that
@@
In the proof we follow the ideas from [26, Proposition 4.6, 4.7]. As $\left|rv_{r}(q)\right|\leq C_{1}$ and $\left|v_{r}(q)-v_{r}(0)\right|\leq C_{2}\bar{q}=C^{\prime}_{2},\,\forall q\in\Omega^{Q}$ by Lemma 38, we can consider a sequence $(r_{n})_{n\in\mathbb{N}}$ converging towards 00 such that the sequences $\big{(}r_{n}v_{r_{n}}(q)\big{)}_{n\in\mathbb{N}}$ and $\big{(}v_{r_{n}}(q)-v_{r_{n}}(0)\big{)}_{n\in\mathbb{N}}$ are convergent for $q\in\Omega^{Q}$. Let $\hat{\gamma}(q)$ denote the limit of the sequence $\big{(}r_{n}v_{r_{n}}(q)\big{)}_{n\in\mathbb{N}}$, we have
@@
Let $z(q)=\hat{w}(q)-\hat{v}(q)$. Since the domain for $z(q)$ is bounded, we know that the supremum and infimum exist. Let us denote $\bar{z}=\sup_{q\in\Omega^{Q}}z(q)$, $\underline{z}=\inf_{q\in\Omega^{Q}}z(q)$ and $\varepsilon=\frac{\hat{\gamma}-\eta}{\bar{z}-\underline{z}+1}$.
Let us first assume $\hat{\gamma}>\eta$ and prove that $\hat{\gamma}\leq\eta$ by contradiction. By the definition of $\varepsilon$, we have, for $\forall q\in\Omega^{Q}$,
$\begin{split}0\leq\varepsilon&(\bar{z}-z(q)+1)\leq\hat{\gamma}-\eta\\ &=H\Big{(}q,(\hat{v}(q^{\prime})-\hat{v}(q))_{q^{\prime}\in\{q-1,q+1\}}\Big{)}-H\Big{(}q,(\hat{w}(q^{\prime})-\hat{w}(q))_{q^{\prime}\in\{q-1,q+1\}}\Big{)}\,.\end{split}$
@@
$\begin{split}-\varepsilon\hat{w}(q)+H\Big{(}q,(\hat{w}(q^{\prime})-&\hat{w}(q))_{q^{\prime}\in\{q-1,q+1\}}\Big{)}\leq\\ &-\varepsilon(\hat{v}(q)+\bar{z}+1)+H\Big{(}q,(\hat{v}(q^{\prime})-\hat{v}(q))_{q^{\prime}\in\{q-1,q+1\}}\Big{)}\,.\end{split}$
By using the comparison principle, see [26, Lemma 4.4], we know that $\hat{v}(q)+\bar{z}+1\leq\hat{w}(q)$, for $\forall q\in\Omega^{Q}$, which indicates a contradiction with the definition of $\bar{z}$. Hence, we have $\hat{\gamma}\leq\eta$. By simply changing the order of $\hat{\gamma}$ and $\eta$, we can obtain that $\hat{\gamma}\geq\eta$. Therefore, we conclude that $\hat{\gamma}=\eta$, i.e. $\hat{\gamma}$ is independent of the sequence $(r_{n})_{n\in\mathbb{N}}$. ∎